# UnrealZoo: Enriching Photo-realistic Virtual Worlds for Embodied AI

### v1

@@ 4 Experiments
In this section, we use a subset of UnrealZoo to demonstrate the usability of the collected environments. For visual navigation, we select two scenes with complex spatial structures to train and validate the RL-based and VLM-based agents. For active tracking, we select at most 8 scenes as training environments and validate the generalization of the learned policy in another 24 scenes, which are divided into four categories according to the scene types. The results demonstrate the importance of the diversity of the training environments to the cross-domain generalization. For social tracking, we analyze the robustness of the agent in social environments with different control frequencies, using the toolkit provided in the gym to generate crowds with varying populations and control frequencies.
4.1 Visual Navigation
visual navigation in the wild introduces a new level of complexity compared to traditional navigation tasks for indoor scenes or autonomous driving, which often run on complex 3D spatial structures. Differently, we place the agent in open-world environments where it must take a set of locomotion, e.g., running, climbing, jumping, crouching, to go over the various obstacles in unstructured terrains to reach the target object. In this setting, the agent requires advanced spatial reasoning and actions to make real-time decisions about its path. The emphasis on such complex environments ensures the agent can operate effectively in a broad range of challenging scenarios, moving beyond the constraints of traditional navigation frameworks. The details of the task setting are introduced in Appendix B.1.
@@ 4.1 Visual Navigation
Evaluation Metrics. We employ two key metrics to evaluate visual navigation agents: 1) Average Episode Length (EL), representing the average number of steps per episode over 50 episodes. 2) Success Rate (SR), measuring the percentage of episodes the agent successfully navigates to the target object out of 50 total episodes, which represents the navigation capability in the wild environment.
@@
Baselines for Navigation. We build simple baselines to demonstrate the applicability of our environments for training reinforcement learning agents and benchmark the agents based on pre-trained large models. 1) Online RL: We trained the RL-based navigation agents separately in the Roof and Factory environments using a distributed online reinforcement learning (RL) approach, e.g. A3C (Mnih et al., 2016). The training curve is shown in Figure 16. The model takes the first-person view segmentation mask and the relative position between the agent and target as input, and outputs direct control signals (from the predefined action space) to navigate. This setup allows the agent to learn and optimize navigation strategies during continuous interaction with the environment. Please refer to Appendix C.1 for the implementation details. 2) GPT-4o: We employ the GPT-4o model to take action, leveraging its powerful multi-modal reasoning capabilities. The model takes first-person view images and the relative position between the agent and the fixed target as input. The GPT-4o model follows our prompt template (See Table 14) as guidance, reasoning appropriate actions from the predefined control space to guide the agent toward the target. 3) Human: We also have a human player control the agent using a keyboard, similar to a first-person video game. The player navigates the agent from a random starting point to a fixed target, making decisions based on visual observations from the shared control space.
@@
Results. In Table 3, we report the performances of different methods in two unstructured scenes. The RL-based agent performs moderately well, achieving better results in the simpler environment (IndustrialArea) compared to the Roof, where the target object is located on different levels of stairs. The GPT-4o agent struggles in both scenarios. This infers that the GPT-4o performs poorly in complex 3D scene reasoning. As a reference, the human player completes both tasks with the fewest steps and a 1.00 success rate, underscoring the significant performance gap between current embodied AI agents and humans, indicating substantial room for improvement to navigate in such complex, open-world environments.
Table 3: The results (EL/SR) of visual navigation in two unstructured terrains.
Methods Roof IndustrialArea
Online RL 1660/0.32 261/0.52
GPT-4o 2000/0.00 369/0.20
Human 515/1.00 158/1.00
@@
Figure 4: An exemplar sequence from the embodied navigation agent in the Roof. The RL-based agent learned to climb on a box and wall and jump over an obstacle to reach the goal location in a short path.
4.2 Active Visual Tracking
We evaluate the generalization of the tracking agents across four environment categories: Interior Scenes, Palaces, Wilds, and Modern Scenes. Each category contains 4 individual environments, as shown in Figure  9. We aim to capture a broad range of features in our environment collection by selecting four distinct and representative scenes from each category, ensuring a comprehensive evaluation of the agents’ capabilities. The details of the tasks are introduced in Appendix B.2. We analyzed the effectiveness of the diversity of the training data by collecting demonstrations with different numbers of training environments.
@@ 4.2 Active Visual Tracking
Evaluation Metrics. Our evaluation employs three key metrics: (1) Average Episodic Return (ER), which calculates the mean episodic return over 50 episodes, providing insights into overall tracking performance; (2) Average Episode Length (EL), representing the average number of steps per episode, which reflects long-term tracking effectiveness; and (3) Success Rate (SR), measuring the percentage of episodes that complete 500 steps out of 50 total episodes.
Baselines for Active Visual Tracking. For the RL-based agents, we extend from the official implementation settings from the recent offline RL method (Zhong et al., 2024), collecting offline datasets and employing the original network architecture. To demonstrate the impact of data diversity on tracking performance, we collect three sets of offline datasets, each containing 100k steps. The key difference between these datasets is the number of environments used for data collection: one was collected in a single environment (denoted as 1 Env.), another in two environments(denoted as 2 Envs.), and the third in eight distinct environments (denoted as 8 Envs.). The offline training curve of each setting is shown in Figure 15. The environment distribution of each dataset setting is shown in Figure 11. It is worth noting that FlexibleRoom, one of the environments used for data collection, is a unique abstract environment, with all objects represented as geometric shapes covered by randomized patterns. This distinctive setup contrasts with the more realistic and diverse environments in the collection, offering a unique scenario for testing agent adaptability. For the VLM-based agents, we utilize the latest large models GPT-4o to directly generate actions based on observed images for tracking a target person. To ensure smooth and precise transitions, we designed a system prompt that helps the model understand the task while standardizing the output format to align with predefined action settings. This prompt ensures the model produces actions coherent with the task’s requirements. Specifically, GPT-4o is tasked with generating concrete action decisions from a predefined instruction space: moving forward, moving backward, turning left, turning right, or maintaining the current position. Once an instruction is generated, we map it to corresponding linear and angular velocities to update the agent’s movement in the environment. It is important to note that while the system prompt can use raw image observations as input, our experience shows poor alignment performance and significant time delays, which pose challenges for real-time tracking. The full system prompt and mapping relationship are provided in Appendix C.2.
@@
Result Analysis. We first evaluate the performance of agents trained with offline datasets collected from varying numbers of environments (1 Env., 2 Envs., 8 Envs.) across 16 distinct environments. We list the detailed evaluation results across the entire 16 environments in Table 11. To better visualize the performance change of different training settings within various scene categories, we calculate the average success rate (SR) of each agent in four categories, the results are shown in Figure 5. The results reveal a clear trend: as the number of environments used for training increases, agent long-term tracking performance generally improves across all categories. In the Wilds, a significant increase in success rate is observed with the 8 Envs. dataset, which involves the highest diversity of environments. This demonstrates that diverse environmental exposure plays a crucial role in improving the agent’s generalization capabilities in more complex, open-world environments. The lower success rate in the 1 Env. dataset highlights the limitations of training solely in abstract settings like the FlexibleRoom. Similarly, in the Palace, the success rate improves notably from 1 Env. to 8 Envs., suggesting that training with a broader range of environments helps the agent better adapt to intricate spatial structures typical of Palace-like maze environments.
@@
Figure 5: Average success rate of agents across four environment categories: Compact Interior, Wildscape Realm, Palace Maze, and Lifelike Urbanity, evaluated under three offline dataset settings (1 Env, 2 Envs., 8 Envs.). The results show the generalization capability improves significantly as more diverse environments are included in the dataset. However, environments with complex spatial structures, such as Compact Interior and Palace Maze, exhibit lower success rates, highlighting challenges in obstacle avoidance and navigation.
4.3 Social Tracking
We further evaluate the tracking agents in a social scenario, where the agent needs to follow the target in crowds. Such a setting contains varying high-dynamics objects with similar appearances, e.g., pedestrians. We can directly apply the population control wrapper in the Gym toolkit to extend the environment used for active tracking to this setting.
Robustness to Active Distractions. A key challenge in active visual tracking tasks is managing active distractions, a critical issue for real-world deployment in crowds. Thus, we conducted an experiment in the DowntownWest and generated crowds with varying numbers of human characters as distractors notated as 4D, 8D, and 10D. The number indicates the number of distractors. We compared the performance of the offline RL method, trained under three dataset configurations (1 Env., 2 Envs., 8 Envs.), against the VLM-based method, evaluating the agents’ ability to maintain robust tracking under these different levels of active distractions.
The results in Table 5 show clear performance differences between the offline RL methods (1 Env., 2 Envs., 8 Envs.) and the GPT-4o model in handling active distractions. As the number of distractors increases, the offline RL methods maintain relatively stable success rates (SR), with the highest performance seen in the 8 Envs. setting, which achieves an SR of 0.8 in the 4D condition and remains robust with slight declines in the 8D and 10D conditions (0.72 and 0.68, respectively). This suggests that the agent benefits from the richer diversity of training data, enabling it to handle increasingly complex crowd scenarios more effectively. On the other hand, the GPT-4o model consistently struggles with active distractions, showing significantly lower average returns (ER) and success rates across all settings. The model’s inability to cope with dynamic, crowded environments is evidenced by its poor performance, particularly in the 10D condition where it records a success rate of just 0.1. This highlights a major limitation of the VLM-based method in dynamic environments with active distractions, as it lacks the temporal consistency and real-time adaptability required for effective tracking. Further evaluation results across different environments are provided in the Appendix D.3.
Cross-Embodiment Generalization. We transfer the agent trained for the human character to the robot dog, which observes the world from a lower perspective. We can see that the results in Table 5 drop, particularly the success rate, indicating that the research community should pay more attention to the cross-embodiment generalization.
@@ 4.3 Social Tracking
The Impact of Control Frequency. We employ the time dilation wrapper to simulate different control frequencies during deployment. The frequency of the perception-control loop is crucial for handling dynamic environments. As is shown in Table 4, when the rate drops below 10 FPS, performance significantly declines. We observe that higher control frequencies enable RL-based agents to perform better in social tracking. These results emphasize the importance of building efficient models for embodied agents, to accomplish tasks in dynamic open worlds.
@@
Table 4: The impact of control frequency on tracking performance. We evaluate the agent (Offline RL 1 Env.) in the FlexibleRoom environment using the time dilation wrapper to simulate varying control frequencies.
ER/ EL/ SR.
3 FPS 184/377/0.34
10 FPS 303/449/0.62
30 FPS 368/482/0.92
w/o Control 275/425/0.74
@@
Table 5: Performance comparison of different methods in the DowntownWest environment with varying numbers of distractors (4D, 8D, 10D). Each cell presents three metrics from left to right: Average Episodic Return (ER), Average Episode Length (EL), and Success Rate (SR).
Method
4D
8D
10D
Offline RL 1 Env. 251/450/0.70 201/406/0.58 230/247/0.64
Offline RL 2 Envs. 309/456/0.74 259/424/0.68 258/428/0.68
Offline RL 8 Envs. 245/458/0.80 225/435/0.72 218/444/0.68
Offline RL 8 Envs. (Robot dog) 220/409/0.48 189/386/0.42 143/367/0.40
GPT-4o -102/264/0.16 -64/270/0.14 -80/240/0.10
4.4 Limitation Analysis and Summary
The current RL method shows some capacity to learn spatial-temporal information and dynamically respond to target movement in most scenarios, but it struggles with executing advanced actions like bypassing obstacles. In compact Interior categories and some special environments such as TerrainDemo, IndustrialArea, and ModularSciFiSeason1, which feature irregular landscapes, narrow passageways, and maze-like structures, the agent often collides with casually placed low-level objects. While the agent can track targets, it’s insufficient to handle unpredictable hindrances, especially in key moments like bypassing corners or tight spaces, which increases the likelihood of failure. This highlights a significant limitation: although the agent can learn and react to its environment, it lacks the higher-level reasoning to anticipate and avoid obstacles effectively. Advanced behaviors like bypassing obstacles are crucial for improving performance, especially in cluttered environments where basic reactive controls are insufficient. Incorporating such reasoning mechanisms would help reduce failure rates, particularly in critical scenarios, and improve overall tracking performance.
For the VLM-based method, one key factor contributing to GPT-4o’s notably poor performance, especially in comparison to the RL methods, is its susceptibility to time delays. From our experience, this issue becomes particularly evident when the target makes abrupt movements, such as turning around. Due to the API’s response lag, the GPT-4o agent struggles to track the target in real time, often losing it before receiving updated instructions. This limitation highlights the difficulty of real-time processing in embodied tracking tasks using models that rely on slower external API communications, underscoring the need for more efficient integration methods for such systems.
@@ 5 Conclusions
In conclusion, we present UnrealZoo, a versatile platform designed to advance embodied AI research. The diversity and complexity of the collected realistic environments challenge agents with varying embodied interaction tasks, such as visual navigation, active tracking across various environments, and social tracking in crowds. The enhanced UnrealCV+ API supports efficient data collection and task customization, enabling seamless interaction for both single and multi-agent systems. These features will open up potential applications like developing spatial intelligence in the 3D world and social intelligence in human-AI society, making our platform a valuable tool for pushing the boundaries of embodied AI in real-world scenarios. Looking forward, we will continue to enrich the virtual worlds with diverse scenes, bodies, and interaction tasks, advancing agents from the virtual realm to reality for a harmonious human-AI society.
Limitations. While our proposed environment provides diverse and complex scenarios for visual navigation, tracking, and other visual-based tasks, it currently lacks high-fidelity physical simulation, limiting the agent’s ability to manipulate objects. Additionally, transferring learned behaviors to different embodied agents poses a challenge, as adapting models to various physical structures and control schemes is not yet seamless. These issues highlight areas for further research to enhance interaction dynamics and improve generalization across diverse agent embodiments.
Acknowledgements
This work was supported by the National Science and Technology Major Project (2022ZD0114904), NSFC-6247070125, NSFC-62406010, the State Key Lab of General Artificial Intelligence at Peking University, and Qualcomm University Research Grant. We express our gratitude to Weichao Qiu for his invaluable support in optimizing UnrealCV and for the insightful discussions. We are also thankful to Tingyun Yan for providing the early version of the human entity and to Jingzhe Lin for his assistance in producing demo videos. Our appreciation extends to the active developers in the open-source community for ensuring UnrealCV’s compatibility with the latest Unreal Engine versions, such as UE 5.4. Furthermore, we thank the contributors in the Unreal Engine Marketplace for offering high-quality content and plugins, which have been instrumental in the development of UnrealZoo.

### v2

@@ 4 Experiments
In this section, we conduct experiments involving a series of challenging tasks on UnrealZoo to show its unique features regarding usability, diversity, efficiency, and extensibility. We begin with a visual navigation task to indicate the suitability of UnrealZoo for distributed online reinforcement learning. This task also highlights how the complex spatial structures within our environments introduce new challenges that the off-the-shelf methods struggle to address effectively. Next, we evaluate agents through active visual tracking to validate the necessity of scaling environmental diversity for improving cross-scene generalization. Finally, our social tracking experiments illustrate the exceptional flexibility of UnrealZoo. By enabling easy modification of simulation control frequencies, population scale, and agent morphologies, UnrealZoo reveals previously overlooked challenges in high-dynamic environments and cross-body generalization.
4.1 Emergent Challenges in Visual Navigation
Taking the visual navigation task as an example, a fundamental skill for embodied agents, we demonstrate the usability of our environments for various approaches, including distributed online reinforcement learning and large vision-language models. Additionally, we highlight the emergent challenges posed by unstructured open-world environments. While traditional approaches often operate within well-defined 3D spatial structures, our environments require agents to navigate through complex, unconstrained terrains. Therefore, agents must execute diverse locomotion skills, including running, climbing, jumping, and crouching, to overcome various environmental obstacles while pursuing target objects. This setting demands sophisticated 3D spatial reasoning capabilities and action selection mechanisms for effective real-time decision-making. The details of the task setting, implementation, and training curve are introduced in the supplementary materials.
@@ 4.1 Emergent Challenges in Visual Navigation
Evaluation Metrics. We employ two key metrics to evaluate visual navigation agents: 1) Average Episode Length (EL), representing the average number of steps per episode over 50 episodes. 2) Success Rate (SR), measuring the percentage of episodes the agent successfully navigates to the target object out of 50 total episodes, which represents the navigation capability in the wild environment. 3) Success weighted by Path Length (SPL) [1], a standard metric that accounts for both success and path efficiency.
@@
Baselines for Navigation. We build simple baselines to demonstrate the applicability of our environments for training reinforcement learning agents and benchmark the agents based on pre-trained large models. 1) Online RL: We train RL-based navigation agents independently in the Roof and Factory environments using a distributed online reinforcement learning (RL) approach, e.g., A3C [27]. The model takes the first-person view segmentation mask and the relative position between the agent and target as input, and outputs discrete actions to navigate. This setup allows the agent to iteratively refine navigation strategies through continuous environmental interaction. 2) GPT-4o: We leverage the multimodal reasoning capabilities of GPT-4o for navigation model. Following a structured prompt template, GPT-4o performs reasoning to select appropriate actions from the action space to guide the agent toward the target location. 3) Human: As a reference, we conduct experiments where human players control agents via keyboard inputs, analogous to first-person video game interactions. The human players use the same action space available to other baselines.
@@
Results. In Table 3, we report the performances of different methods in two unstructured scenes. The RL-based agent performs moderately well, achieving better results in the simpler environment (Industrial Area) compared to the Roof, where the target object is located on different levels of stairs. The GPT-4o agent struggles in both scenarios. This implies that the GPT-4o performs poorly in complex 3D scene reasoning. As a reference, the human player completes both tasks with the fewest steps and a 1.00 success rate, underscoring the significant performance gap between current embodied AI agents and humans, indicating substantial room for improvement to navigate in such complex, open-world environments.
Table 3: The results (EL/SR/SPL) of visual navigation agents in two unstructured terrains.
Online RL GPT-4o Human Roof 1660/0.32/0.10 2000/0.00/0.00 515/1.00/0.97 Industrial Area 261/0.52/0.31 369/0.20/0.08 158/1.00/0.98
@@
Figure 5: An exemplar sequence from the embodied navigation agent in the Roof. The RL agent learned to climb on a box and wall and jump over an obstacle to reach the goal in a short path.
4.2 Scaling the Diversity for Active Visual Tracking
In this section, we aim to analyze the effectiveness of the diversity of the training data on the generalization of the active visual tracking. We collect 3 sets of demonstrations from different numbers of environments in UnrealZoo and train the tracking policy via offline reinforcement learning [49]. We evaluate the tracking performance across 16 unseen environments covering four categories: Interior Scenes, Palaces, Wilds, and Modern Scenes. Each category contains 4 individual environments. The implementation details are introduced in the supplementary materials.
@@ 4.2 Scaling the Diversity for Active Visual Tracking
Evaluation Metrics. Our evaluation employs three key metrics: (1) Average Episodic Return (ER): calculates the mean episodic return over 50 episodes, providing insights into overall tracking performance; (2) Average Episode Length (EL): represents the average number of steps per episode, which reflects long-term tracking effectiveness; and (3) Success Rate (SR): measures the percentage of episodes that complete 500 steps out of 50 total episodes.
Baselines for Active Visual Tracking. 1) PID: A classical control approach that uses a PID controller to adjust agent actions by maximizing IoU between the detected target bounding box and an expected position. 2) RL: We adapt the offline reinforcement learning implementation from Zhong et al. [49], using their network architecture while collecting three distinct offline datasets (100k steps each) with varying environmental diversity: single environment (1 Env.), two environments (2 Envs.), and eight environments (8 Envs.). Notably, FlexibleRoom represents an abstract environment where objects appear as geometric shapes with randomized patterns used in the three settings. 3) OpenVLA: We fine-tune OpenVLA [19], an end-to-end large model originally designed for robotic manipulation with strong generalization capabilities, on our most diverse dataset (8 Envs.) to adapt it for tracking tasks. 4) GPT-4o: We implement a VLM-based agent using GPT-4o to generate actions directly from observed images. We designed a prompt template to standardize outputs into predefined action commands (moving forward/backward, turning left/right, or maintaining position), which are then mapped to appropriate velocities. Despite GPT-4o’s capabilities, we observed poor alignment performance and significant latency issues when processing raw image inputs, limiting its effectiveness for real-time tracking.
@@
Result Analysis. We first evaluate the performance of agents trained with offline datasets collected from varying numbers of environments (1 Env., 2 Envs., 8 Envs.) across 16 unseen environmentsTo visualize the performance change of different settings within various scene categories, we calculate the average success rate (SR) of each agent in four categories, and the results are shown in Figure 6. Then we compared the baseline methods’ performance, as shown in Table 5 (0D). The results reveal a clear trend: as the number of environments used for training increases, agent long-term tracking performance generally improves across all categories. In the Wilds, a significant increase in success rate is observed with the 8 Envs. dataset, which involves the highest diversity of environments. This demonstrates that diverse environmental exposure plays a crucial role in improving the agent’s generalization capabilities in more complex, open-world environments. The lower success rate in the 1 Env. highlights the limitations of training solely in abstract settings like the FlexibleRoom. Similarly, in the Palace, the success rate improves notably from 1 Env. to 8 Envs., suggesting that training with a broader range of environments helps the agent better adapt to intricate 3D spatial structures.
@@
Figure 6: Average success rates across four environment categories: Interior Scenes, Palaces, Wilds, and Modern Urbanity. Results compare Offline RL agents trained on three offline dataset settings with increasing environmental diversity (1 Env., 2 Envs., 8 Envs.). Performance consistently improves with dataset diversity, demonstrating enhanced generalization capabilities. Notably, environments with complex spatial structures (Interior Scenes and Palaces) present greater challenges for navigation and obstacle avoidance, resulting in lower success rates.
4.3 Benchmarking Agents at Variant Settings
We augment the environment employed for active tracking by seamlessly integrating the Gym toolkit into diverse settings. We use the population control wrapper to create social tracking scenarios. In this setting, the agent is required to follow the target through crowds. We can also adjust the morphology of the trackers for evaluating the cross-embodiment generalization. Moreover, we employ the time dilation wrapper to simulate a variety of control frequencies. This allow us to analyze the impact of the latency within the perception-control loop on tracking performance.
Robustness to Active Distractions. A key challenge in active visual tracking is handling visual distractions and dynamic obstacles, which is crucial for real-world deployment in crowds. We evaluated agents in four representative environments: Storage House, Desert Ruins, TerrainDemo, and SuburbNeighborhoodDay, with varying numbers of human distractors (0D, 4D, 10D) and compared RL-based methods (trained in 8 environments, denoted as RL 8 Envs.) against PID, OpenVLA, and GPT-4o. Table 5 shows that RL-based methods maintain stable success rates, RL 8 Envs. achieving 0.76 SR in 0D, 0.68 in 4D and 0.56 in 10D, benefiting from diverse training data. In contrast, PID, OpenVLA, and GPT-4o struggle, showing significantly lower returns and success rates. PID suffers from delayed reactions when the target is occluded, leading to failures. OpenVLA tracks effectively at short distances but loses accuracy over time due to accumulated errors. Both OpenVLA and GPT-4o are highly prone to distractions, often tracking the wrong target. These results highlight the limitations of large models in dynamic environments with active distractions.
Cross-Embodiment Generalization. To investigate whether policies can effectively transfer to agents with different morphologies, we apply the agents (RL and OpenVLA) trained on a human character to a quadrupedal robot dog. This exploration is motivated by the practical necessity in real-world applications where a single policy might need to operate across various embodiments with different perceptual perspectives and movement capabilities. As shown in Table  5, the performance decreases notably, particularly the success rate (from 0.68 to 0.48 with 4D and from 0.56 to 0.40 with 10D for RL 8 Envs.), indicating that the community should pay more attention to the cross-embodiment generalization for real-world deployment.
@@ 4.3 Benchmarking Agents at Variant Settings
The Impact of Control Frequency. In high-dynamic scenes, the latency of the perception-control loop plays a critical role in managing multiple dynamic factors effectively. To verify this, We employ the time dilation wrapper to simulate different control frequencies during deployment. As shown in Table 4, when the rate drops below 10 FPS, performance significantly declines. Our results show that RL-based agents achieve superior performance in social tracking scenarios when operating at higher control frequencies. These findings emphasize the necessity of developing efficient models for embodied agents to successfully execute tasks in dynamic social worlds.
@@
Table 4: The impact of control frequency on tracking performance. We evaluate the RL agent (1 Env.) in the FlexibleRoom using the time dilation wrapper to simulate varying control frequencies.
3 FPS 10 FPS 30 FPS w/o FPS Control
ER 184 303 368 275
EL 377 449 482 425
SR 0.34 0.62 0.92 0.74
@@ 3 FPS 10 FPS 30 FPS w/o FPS Control
Table 5: We evaluate different methods in four representative environments: Storage House, Desert Ruins, TerrainDemo, and SuburbNeighborhoodDay, with varying numbers of distractors (0D, 4D, 10D). Each cell presents the averaged values across four environments for three metrics: Episodic Return (ER), Episode Length (EL), and Success Rate (SR). OpenVLA and RL are trained on the dataset collected from 8 environments.
Method 0D 4D 10D PID 113/298/0.34 107/295/0.30 78/105/0.14 GPT-4o 2/228/0.28 -102/264/0.16 -80/240/0.10 OpenVLA 55/323/0.10 -39/149/0.02 -30/156/0.00 RL 321/468/0.76 280/437/0.68 187/402/0.56 PID (Robot dog) 152/245/0.30 149/237/0.22 100/102/0.16 GPT-4o (Robot dog) -5/263/0.20 -98/272/0.16 -59/223/0.04 OpenVLA(Robot dog) 42/301/0.08 -37/142/0.00 -13/117/0.00 RL (Robot dog) 279/433/0.60 220/409/0.48 143/367/0.40
@@ 5 Conclusions
In conclusion, UnrealZoo addresses critical gaps in embodied AI research by providing unprecedented environmental diversity across 100 scenes and 67 embodiment options. By offering easy-to-use Python APIs and toolkits, we support efficient environmental interactions, distributed training, and task customization in UnrealZoo. The experiments on UnrealZoo reveal the importance of environment diversity for agent generalization capabilities, the necessity of low-latency control in dynamic scenes, and key challenges faced by current methods for embodied agents in open worlds. These findings highlight a clear path forward for developing more robust embodied AI systems capable of navigating the complexity of real-world deployment.
Looking forward, UnrealZoo will drive research in several key areas: 1) spatial intelligence development in complex 3D environments; 2) social intelligence research through multi-agent interactions; 3) safe pre-deployment evaluation and limitation identification; 4) cross-embodiment learning and skill transfer; and 5) adaptation to dynamic environments. As we continue to enrich virtual worlds and interaction capabilities, UnrealZoo will play an increasingly vital role in advancing agents from simulation to reality, ultimately fostering harmonious human-AI coexistence.
Acknowledgments
This work was supported by the National Science and Technology Major Project (2022ZD0114904), NSFC-6247070125, NSFC-62406010, the Funding Scheme for Research and Innovation of FDCT (Grant No. 0021/2025/ITP1), the Fundamental Research Funds for the Central Universities, and Qualcomm University Research Grant. We express our gratitude to Weichao Qiu for his invaluable support in optimizing UnrealCV and for the insightful discussions. We are also thankful to Tingyun Yan for providing the early version of the human entity and to Jingzhe Lin for his assistance in producing demo videos. Our appreciation extends to the active developers in the open-source community for ensuring UnrealCV’s compatibility with the latest Unreal Engine versions, such as UE 5.4. Furthermore, we thank the contributors in the Unreal Engine Marketplace for offering high-quality content and plugins, which have been instrumental in the development of UnrealZoo.